# MA-GTS: A Multi-Agent Framework for Solving Complex Graph Problems in Real-World Applications

### v1

@@ Abstract
Graph-theoretic problems arise in real-world applications like logistics, communication networks, and traffic optimization. These problems are often complex, noisy, and irregular, posing challenges for traditional algorithms. Large language models (LLMs) offer potential solutions but face challenges, including limited accuracy and input length constraints. To address these challenges, we propose MA-GTS (Multi-Agent Graph Theory Solver), a multi-agent framework that decomposes these complex problems through agent collaboration. MA-GTS maps the implicitly expressed text-based graph data into clear, structured graph representations and dynamically selects the most suitable algorithm based on problem constraints and graph structure scale. This approach ensures that the solution process remains efficient and the resulting reasoning path is interpretable. We validate MA-GTS using the G-REAL dataset, a real-world-inspired graph theory dataset we created. Experimental results show that MA-GTS outperforms state-of-the-art approaches in terms of efficiency, accuracy, and scalability, with strong results across multiple benchmarks (G-REAL 94.2%, GraCoRe 96.9%, NLGraph 98.4%). MA-GTS is open-sourced at https://github.com/ZIKEYUAN/MA-GTS.git.
@@
Zike Yuan1,2, Ming Liu1, Hui Wang2, Bing Qin1 1Harbin Institute of Technology, Shenzhen, China, 2Peng Cheng Laboratory, Shenzhen, China {yuanzk,wangh06}@pcl.ac.cn {mliu,qinb}@ir.hit.edu.cn
Figure 1: This figure presents MA-GTS framework for solving real-world graph problems, consisting of three layers: Information Extraction, Knowledge Integration, and Algorithm Execution, each with specialized agents to process and solve the problem.
@@ 1 Introduction
Graph-theoretic problems have extensive applications in domains such as logistics scheduling, communication networks, production planning, and traffic optimizationLi et al. (2023b). These problems typically involve a large number of nodes and edges, coupled with complex constraints and dynamic variations, making their solution highly challengingBondy and Murty (2008). Despite significant advancements in graph theory and algorithmic design, traditional approaches remain computationally expensive and inefficient when handling large-scale, high-complexity problems. Existing methods, including exact algorithms, greedy strategies, and dynamic programmingBellman (1966), perform well on small-scale instances. However, as problem size increases, their computational complexity and memory requirements grow exponentially, rendering them impractical for real-world applications. While heuristic methodsKokash (2005) can improve performance under specific conditions, they often suffer from local optima and require extensive parameter tuning and model selection. Therefore, developing efficient and scalable solution frameworks capable of addressing the computational demands and structural variability of complex graph-theoretic problems remains a critical research challenge.
@@
Recent advancements in large language models (LLMs) have spurred interest in their applications for graph-theoretic problems. Leveraging their natural language processing (NLP) capabilities, LLMs can serve as scene interpreters (mapping real-world problems to graph models), graph extractors (identifying graph structures from unstructured data), and graph algorithm invokers (assisting in solving and optimizing graph-based problems), addressing certain limitations of traditional algorithms. However, significant challenges remain. Firstly, LLMs rely on statistical pattern matching rather than strict mathematical computations, limiting their reasoning accuracy and making them unreliable for NP-hard problemsHochba (1997). Secondly, their ability to handle large-scale graphs is limited by the TransformerVaswani (2017) architecture’s context window and computational complexity, which restricts their capacity to capture global information. Finally, LLMs lack the ability to decompose and map real-world graph theory problems, which often contain complex textual noise and implicit graph structures. As a result, LLMs struggle with accurate denoising and may mismap node information, reducing reasoning interpretability. These limitations highlight the inadequacy of single-agent LLM frameworks for solving complex graph-theoretic problems in real-world applications and underscore the need for more efficient and scalable paradigms.
@@
To tackle these challenges, we propose MA-GTS(Multi-Agent Graph Theory Solver), an innovative multi-agent framework designed to address complex graph-theoretic problems through agent collaboration and competition. Figure 1 illustrates the framework, which incorporates a multi-agent coordination mechanism allowing agents to perform local searches independently while sharing information and cooperating, thus improving solution efficiency and accuracy. MA-GTS analyzes the original real-world problem textual data, filters out noise, and extracts key graph data and problem-specific details, reducing the text length that LLMs must process and enhancing reasoning efficiency. This coordination mitigates the limitations of LLMs in implicit graph structure modeling, ensuring efficient solutions for complex graph tasks. Additionally, dynamic agent interactions enable the framework to address large-scale problems and adapt to complex constraints and dynamic changes.
@@
To validate the effectiveness of the multi-agent framework, we introduce the G-REAL dataset, designed to simulate complex graph theory problems relevant to real-world scenarios. Unlike traditional datasets that rely on simple textual descriptions of graph structures, G-REAL better reflects practical applications for large-scale models. Experiments comparing MA-GTS with state-of-the-art open-source and closed-source LLMs—three closed-source and two open-source models—using both direct and Chain of Thought (CoT) reasoning, show that MA-GTS significantly outperforms existing LLMs in terms of efficiency and accuracy. Notably, it excels in solving large-scale problems with complex constraints, offering superior scalability, robustness, and cost-effectiveness. The primary contributions of this study are as follows:
@@
Figure 2: This figure provides a detailed description of the composition and characteristics of the G-REAL dataset, as well as the entire graph theory problem-solving pipeline of MA-GTS, including specific functions and the formats of input and output.
@@ 2 Related Work
LLMs for Graph: Recent advancements in LLMs for graph tasks have led to significant contributions in methodology and evaluation. These tasks are often classified into Enhancer, Predictor, and Alignment types (Li et al., 2023b). Notably, (Pan et al., 2023) presents a roadmap for unifying LLMs with Knowledge Graphs (KGs), while (Chai et al., 2023) proposes an end-to-end method for solving graph-related problems,Cao et al. (2024) improves LLMs’ understanding of graph structures by addressing positional biases and incorporating an external knowledge base. On the evaluation front, several benchmarks have been introduced. NLGraph (Wang et al., 2024) offers a simple test dataset for graph tasks, and GPT4Graph (Guo et al., 2023) evaluates LLM capabilities on semantic tasks. GraCoReYuan et al. (2025) comprehensively verifies the graph understanding and reasoning capabilities of LLM. Other notable works include (Liu and Wu, 2023), which assesses LLMs in graph data analysis, and (Perozzi et al., 2024), which designs a hint method for graph tasks.
LLM Agents: In recent years, several multi-agent frameworks have been proposed to enhance the coordination and efficiency of language models in complex tasks. MetaGPTHong et al. (2023) reduces hallucinations in complex tasks by embedding human workflows into language models. CAMELLi et al. (2023a) promotes autonomous cooperation among agents, guiding them to align with human goals and studying their interactions. AutoGenWu et al. (2023) is a flexible framework that allows developers to customize agent interactions using both natural language and code, suitable for various fields. In addition, Li et al. (2024a) can be used to solve simple graph problems, Li et al. (2024b) is an autonomous agent that uses LLMs to create animated videos from simple narratives.
@@ 3 MA-GTS
The IEL parses textual and structured data, extracts graph-related information, and identifies problem types to provide standardized inputs. The KIL constructs structured graph data, applying graph theory and optimization strategies to improve accuracy and scalability. The AEL calls specified algorithms for computation and performs self-checking to solve complex graph problems efficiently. Figure 2 illustrates the functionality of each agent in their respective layers.
@@ 3.1 Information Extraction Layer (IEL)
The IEL extracts relevant information from text and unstructured data, converting it into a structured format for downstream processing. It also filters out irrelevant content, refining problem-specific information to enhance LLM reasoning. Additionally, it captures implicit graph-structural information, improving problem-solving efficiency and mitigating the impact of text length on model comprehension and inference.
Textual Information Extraction Agent (TIEA): The TIEA analyzes real-world graph-theoretic problems, extracting key textual information not directly related to the graph structure or solution objectives. Using NLP techniques, it identifies and structures contextual descriptions, background information, entities, concepts, and definitions. Its goal is to organize context, terminology, and related concepts, providing semantic support for subsequent analysis and problem solving. The extracted information is then output in a standardized format for downstream processing.
Graph Structure Information Extraction Agent (GSIEA): The GSIEA extracts implicitly embedded graph-structural information from text, particularly structured formats like tables, lists, adjacency matrices, or edge lists. It parses these inputs to identify nodes, edges, weights, and other topological properties, converting them into standardized graph representations (e.g., adjacency matrices or lists). This transformation enables downstream agents to efficiently use the extracted data for problem solving.
Problem Information Extraction Agent (PIEA): The PIEA leverages LLMs’ problem classification capabilities to analyze real-world graph-theoretic problems, identify their types, and extract key components. It classifies problems (e.g., shortest path, network flow, graph matching), extracts relevant constraints and objectives, and outputs the information in a structured format. This guidance improves the accuracy and efficiency of downstream problem-solving agents.
Formally,the operation of IEL is:
$\mathcal{T}\leftarrow A_{TIEA}(P),$ (1)
$\mathcal{P}\leftarrow A_{PIEA}(P),$
$G\leftarrow A_{GSIEA}(P)$
$IEL_{output}=(\mathcal{T},\mathcal{P},G)$
@@ 3.2 Knowledge Integration Layer (KIL)
Structured Graph Information Agent (SGIA): The SGIA standardizes the graph structure data extracted by the GSIEA, ensuring efficiency, consistency, and usability. It converts raw data into standard formats compatible with various environments, while performing data cleaning, deduplication, and format optimization to maintain accuracy. Additionally, it optimizes data storage and indexing based on algorithm requirements, enhancing computational efficiency for large-scale graphs. Without this agent, data inconsistencies, redundancy, and unoptimized structures could hinder algorithm performance. As a key component of MA-GTS, it ensures data standardization and optimization for efficient, scalable problem-solving.
Graph Theory Agent (GTA): The GTA integrates information from the TIEA and PIEA to analyze graph-theoretic problems and determine optimal solution strategies using a pre-stored Graph Theory Knowledge Base, enhancing LLM inference efficiency. It models the input problem by extracting key features such as type, constraints, and structural complexity, then queries the Graph Theory Knowledge Base to select the most suitable solution method from classical algorithms (e.g., shortest path, maximum flow, graph matching)Gallo and Pallottino (1988); Goldberg and Tarjan (1988) and heuristic techniques. By matching problems to algorithms, it reduces inefficient exhaustive searches, cutting computational costs and improving solution quality. Additionally, it provides guidance for multi-agent collaboration, enabling the AEL to invoke the optimal algorithm directly, ensuring efficiency and scalability. Without this agent, LLMs may face suboptimal strategy selection, excessive computation, and reduced efficiency. As a core component of MA-GTS, it is vital for algorithm selection and inference optimization in complex graph-theoretic tasks.
Formally,the operation of KIL is:
$\mathcal{L}_{\mathcal{P}}\leftarrow A_{GTA}(\mathcal{T},\mathcal{P},\mathcal{K_{G}}),$ (2)
$Alg^{*}\leftarrow\arg\operatorname{opt}_{Alg_{i}\in\mathcal{L}_{\mathcal{P}}}A_{GTA}(Alg_{i},\mathcal{T}),$
$G^{\prime}\leftarrow A_{SGIA}(G),$
$KIL_{output}=(G^{\prime},Alg^{*})$
where $\mathcal{L}_{\mathcal{P}}$ represents the set of graph theory algorithms selected by GTA based on textual and problem-specific information, $\mathcal{K_{G}}$ denotes the Graph Theory Knowledge Base, $Alg^{*}$ refers to the algorithm most suitable for the given graph size, and $G^{\prime}$ stands for the normalized graph structure data.
@@ 3.3 Algorithm Execution Layer (AEL)
Algorithm Solving Agent (ASA): The ASA is the core computational unit of the AEL, responsible for solving problems by executing algorithmic functions based on the optimal strategy selected by the GTA and the structured graph data processed by the SGIA. It utilizes a Graph Theory Algorithm Library that integrates exact algorithmsNoto and Sato (2000) and heuristic approaches, ensuring optimal or near-optimal solutions across various problem scenarios. After computation, the agent performs result integration and verification through cross-validation, error analysis, and constraint checking to ensure correctness. It also provides explainable reasoning, offering inference paths, key decision points, and optimization steps to enhance transparency. As the computational core of MA-GTS, the ASA enables efficient, robust, and scalable solutions for complex graph-theoretic problems.
Formally,the operation of AEL is:
$Code_{Alg^{*}}\leftarrow A_{ASA}(Alg^{*},\mathcal{L}_{code}),$ (3)
$S_{code}\leftarrow A_{ASA}Coding(Code_{Alg^{*}},G^{\prime}),$
$S^{0}\leftarrow A_{ASA}(S_{code},Alg^{*},G^{\prime}),$

### v2

@@ Abstract
Graph-theoretic problems arise in real-world applications like logistics, communication networks, and traffic optimization. These problems are often complex, noisy, and irregular, posing challenges for traditional algorithms. Large language models offer potential solutions but face several challenges, including limited accuracy, input length constraints, and suboptimal algorithm selection. To address these challenges, we propose MA-GTS (Multi-Agent Graph Theory Solver), a multi-agent framework that decomposes these complex problems through agent collaboration. MA-GTS maps the implicitly expressed text-based graph data into clear, structured graph representations and dynamically selects the most suitable algorithm based on problem constraints and graph structure scale. We validate MA-GTS using the G-REAL dataset, a real-world-inspired graph theory dataset we created. Experimental results show that MA-GTS outperforms state-of-the-art methods in cost-effectiveness, accuracy, and scalability, achieving strong results on multiple benchmarks (G-REAL 93.6%, GraCoRe 96.9% NLGraph 98.4%) with robust performance on both closed- and open-source base models. MA-GTS and G-REAL are open-sourced at https://github.com/ZIKEYUAN/MA-GTS.git.
@@
Zike Yuan1,2, Ming Liu1,2,*, Hui Wang2,*, Bing Qin1,2,* 1Harbin Institute of Technology, Shenzhen, China, 2Peng Cheng Laboratory, Shenzhen, China {yuanzk,wangh06}@pcl.ac.cn {mliu,qinb}@ir.hit.edu.cn
footnotetext: * B. Qin, H. Wang and M. Liu are corresponding authors.
@@ 1 Introduction
Graph-theoretic problems have extensive applications in domains such as logistics scheduling, communication networks, production planning, and traffic optimization suvery. These problems typically involve a large number of nodes and edges, coupled with complex constraints and dynamic variations, making their solution highly challenging bondy2008graph. Despite significant advancements in graph theory and algorithmic design, traditional approaches remain computationally expensive and inefficient when handling large-scale, high-complexity problems. Existing methods, including exact algorithms, greedy strategies, and dynamic programming bellman1966dynamic, perform well on small-scale instances. However, as problem size increases, their computational complexity and memory requirements grow exponentially, rendering them impractical for real-world applications. While heuristic methods kokash2005introduction can improve performance under specific conditions, they often suffer from local optima and require extensive parameter tuning and model selection. Therefore, developing efficient and scalable solution frameworks capable of addressing the computational demands and structural variability of complex graph-theoretic problems remains a critical research challenge.
Figure 1: MA-GTS leverages multi-agent collaboration to overcome noise and semantic loss in real-world graph problems, leading to better answers.
Figure 2: MA-GTS framework for solving real-world graph problems, consisting of three layers: Information Extraction, Knowledge Integration, and Algorithm Execution, each with specialized agents.
@@
Recent advancements in LLMs have spurred interest in their applications for graph-theoretic problems. Leveraging their natural language processing (NLP) capabilities, LLMs can serve as scene interpreters (mapping real-world problems to graph models), graph extractors (identifying graph structures from unstructured data), and graph algorithm invokers (assisting in solving and optimizing graph-based problems), addressing certain limitations of traditional algorithms. However, significant challenges remain in existing methods (LLMs and simple multi-agent framework). Figure 1 clearly illustrates the challenges existing methods face when addressing real-world graph problems. Firstly, LLMs rely on statistical pattern matching rather than strict mathematical computations, limiting their reasoning accuracy and making them unreliable for NP-hard problems hochba1997approximation. Secondly, their ability to handle large-scale graphs is limited by the Transformer vaswani2017attention architecture’s context window and computational complexity, which restricts their capacity to capture global information. Finally, LLMs lack the ability to decompose and map real-world graph theory problems, which often contain complex textual noise and implicit graph structures. In summary, existing methods struggle to effectively handle long texts and graph problems in real-world scenarios. Problems like disordered nodes, noisy text, and poor algorithm choices can all affect the quality of graph modeling, text understanding, and the interpretability of reasoning. These limitations highlight the inadequacy of existing methods for solving complex graph-theoretic problems in real-world applications and underscore the need for more efficient and scalable paradigms.
@@
To tackle these challenges, we propose MA-GTS(Multi-Agent Graph Theory Solver), an innovative multi-agent framework designed to address complex real world graph-theoretic problems through agent collaboration and competition. Figure 2 illustrates the framework, which incorporates a multi-agent coordination mechanism allowing agents to perform local searches independently while sharing information and cooperating, thus improving solution efficiency and accuracy. MA-GTS analyzes the original real-world problem textual data, filters out noise, and extracts key graph data and problem-specific details, reducing the text length that LLMs must process and enhancing reasoning efficiency. MA-GTS selects the optimal graph algorithm based on refined text and adjusts the graph’s textual representation to match the algorithm, improving reasoning and solution quality. This coordination mitigates the limitations of LLMs in implicit graph structure modeling, ensuring efficient solutions for complex graph tasks. Additionally, dynamic agent interactions enable the framework to address large-scale problems and adapt to complex constraints and dynamic changes.
Figure 3: This figure details the G-REAL dataset’s composition and features, along with the full MA-GTS graph problem-solving pipeline, outlining each component’s functions and input/output formats.
@@
To validate the effectiveness of the multi-agent framework, we introduce the G-REAL dataset, designed to simulate complex graph theory problems relevant to real-world scenarios. Unlike traditional datasets that rely on simple textual descriptions of graph structures, G-REAL better reflects practical applications for large-scale models. Experiments comparing MA-GTS with state-of-the-art open-source and closed-source LLMs (including three closed-source and three open-source models), as well as with a general multi-agent framework and a graph-specific multi-agent framework, show that MA-GTS significantly outperforms existing LLMs and multi-agent frameworks in terms of efficiency and accuracy, under both direct reasoning and Chain of Thought (CoT) wei2022chain reasoning settings. Notably, it excels in solving large-scale problems with complex constraints, offering superior scalability, robustness, and cost-effectiveness. The primary contributions of this study are as follows:
@@ 2 Related Work
LLMs for Graph: Recent advancements in LLMs for graph tasks have led to significant contributions in methodology and evaluation. These tasks are often classified into Enhancer, Predictor, and Alignment types suvery. Notably, 2 presents a roadmap for unifying LLMs with Knowledge Graphs (KGs), while 3 proposes an end-to-end method for solving graph-related problems, cao2024graphinsight improves LLMs’ understanding of graph structures by addressing positional biases and incorporating an external knowledge base. On the evaluation front, several benchmarks have been introduced. NLGraph NLgraph-6 offers a simple test dataset for graph tasks, and GPT4Graph gpt4graph-7 evaluates LLM capabilities on semantic tasks. GraCoRe yuan2025gracore comprehensively verifies the graph understanding and reasoning capabilities of LLM. In addition to these representative benchmarks, ProGraph li2024can, GraphArena tang2024grapharena, GLBench li2024glbench, etc. are also widely used. Other notable works include (8), which assesses LLMs in graph data analysis, and (10), which designs a hint method for graph tasks.
LLM Agents: Several multi-agent frameworks have been proposed to improve coordination and efficiency in complex tasks. MetaGPT hong2023metagpt embeds human workflows into LLMs to reduce hallucinations. CAMEL li2023camel enables autonomous agent cooperation aligned with human goals, and its extension OWL owl2025 builds on this. AutoGen wu2023autogen offers a flexible framework for customizing agent interactions via natural language and code. Additionally, li2024graphteam addresses simple graph problems. Multi-agent frameworks like GraphTeam li2024graphteam, GCoder zhang2024gcoder, and GraphAgent hu2024scalable can enhance the reasoning ability of LLMs through multiple interactions, but they are mainly applied to standard graph structures, and their effectiveness on real-world graph theory problems remains uncertain.
@@ 3 MA-GTS
We consider a real-world graph problem $P$, modeled as a graph $G$. The system uses a graph theory knowledge base $\mathcal{K}_{\mathcal{G}}$ and an algorithm library $\mathcal{L}_{\text{code}}$ to support problem understanding and solving. From $P$, it extracts the textual description $\mathcal{T}$, identifies the problem type $\mathcal{P}$, and constructs the graph structure $G$. The objective is to automatically select an appropriate algorithm $Alg^{*}$, apply it to the structured graph $G^{\prime}$, and iteratively optimize the solution $S^{n}$.
@@
The IEL processes text and structured data to extract graph information and identify problem types for standardized input. The KIL builds structured graph data using graph theory and optimization to enhance accuracy and scalability. The AEL runs specified algorithms and performs self-checks to efficiently solve complex graph problems. Figure 3 shows each agent’s function by layer.
@@ 3.1 Information Extraction Layer (IEL)
The IEL extracts relevant information from text and unstructured data, structures it for downstream use, and filters out irrelevant content to sharpen problem-specific details. It also captures implicit graph structures to boost efficiency and reduce the effects of text length on LLMs inference.
Textual Information Extraction Agent (TIEA): The TIEA analyzes real-world graph problems to extract key textual information unrelated to graph structure or solution goals. Using NLP, it identifies and structures context, background, entities, concepts, and definitions, organizing semantic content to support later analysis. The output is standardized for downstream processing.
@@
Graph Structure Information Extraction Agent (GSIEA): The GSIEA extracts implicitly embedded graph-structural information from text, particularly structured formats like tables, lists, adjacency matrices, or edge lists. It parses these inputs to identify nodes, edges, weights, and other topological properties, converting them into standardized graph representations (e.g., adjacency matrices, lists). This transformation enables downstream agents to efficiently use the extracted data for problem solving.
Problem Information Extraction Agent (PIEA): The PIEA leverages LLMs’ problem classification capabilities to analyze real-world graph-theoretic problems, identify their types, and extract key components. It classifies problems (e.g., shortest path, network flow, graph matching), extracts relevant constraints and objectives, and outputs the information in a structured format. This guidance improves the accuracy and efficiency of downstream problem-solving agents. Formally,the operation of IEL is:
$\begin{gathered}\mathcal{T}\leftarrow A_{TIEA}(P),\mathcal{P}\leftarrow A_{PIEA}(P),\\ G\leftarrow A_{GSIEA}(P),IEL_{output}=(\mathcal{T},\mathcal{P},G)\end{gathered}$ (1)
@@ 3.2 Knowledge Integration Layer (KIL)
Structured Graph Information Agent (SGIA): The SGIA standardizes graph data from the GSIEA for efficient, consistent, and usable output. It cleans, deduplicates, and optimizes raw data into formats compatible with diverse environments to ensure accuracy. Additionally, it optimizes data storage and indexing based on algorithm requirements, enhancing computational efficiency for large-scale graphs. Without this agent, data inconsistencies, redundancy, and unoptimized structures could hinder algorithm performance. As a key component of MA-GTS, it ensures data standardization and optimization for efficient, scalable problem-solving.
Graph Theory Agent (GTA): The GTA integrates information from the TIEA and PIEA with a Graph Theory Knowledge Base to analyze graph problems and find optimal solutions, improving LLM inference efficiency. It models the input problem by extracting key features such as type, constraints, and structural complexity, then queries the Graph Theory Knowledge Base to select the most suitable solution method from classical algorithms (e.g., shortest path, maximum flow, graph matching) gallo1988shortest; goldberg1988new and heuristic techniques. By matching problems to algorithms, it reduces inefficient exhaustive searches, cutting computational costs and improving solution quality. Additionally, it also guides multi-agent collaboration, allowing the AEL to directly invoke optimal algorithms for efficient, scalable execution. Without it, LLMs risk poor strategy selection, high computation, and lower efficiency. As a key MA-GTS component, it ensures effective algorithm selection and inference in complex graph tasks. Formally,the operation of KIL is:
$\begin{gathered}\mathcal{L}_{\mathcal{P}}\leftarrow A_{GTA}(\mathcal{T},\mathcal{P},\mathcal{K_{G}}),\\ Alg^{*}\leftarrow\arg\operatorname{opt}_{Alg_{i}\in\mathcal{L}_{\mathcal{P}}}A_{GTA}(Alg_{i},\mathcal{T}),\\ G^{\prime}\leftarrow A_{SGIA}(G),KIL_{output}=(G^{\prime},Alg^{*})\end{gathered}$ (2)
where $\mathcal{L}_{\mathcal{P}}$ represents the set of graph theory algorithms selected by GTA based on textual and problem-specific information, $\mathcal{K_{G}}$ denotes the Graph Theory Knowledge Base, $Alg^{*}$ refers to the algorithm suitable for the given graph size, and $G^{\prime}$ stands for the normalized graph structure data.
@@ 3.3 Algorithm Execution Layer (AEL)
Algorithm Solving Agent (ASA): The ASA is the core computational unit of the AEL, responsible for solving problems by executing algorithmic functions based on the optimal strategy selected by the GTA and the structured graph data processed by the SGIA. It utilizes a Graph Theory Algorithm Library that integrates exact algorithms noto2000method and heuristic approaches, ensuring suitable solutions across various problem scenarios. After computation, the agent performs result integration and verification through cross-validation, error analysis, and constraint checking to ensure correctness. The ASA also offers explainable reasoning with inference paths, key decisions, and optimization steps for transparency. As MA-GTS’s computational core, it delivers efficient, robust, and scalable solutions for complex graph problems. Formally,the operation of AEL is:
$\begin{gathered}Code_{Alg^{*}}\leftarrow A_{ASA}(Alg^{*},\mathcal{L}_{code}),\\ S_{code}\leftarrow A_{ASA}Coding(Code_{Alg^{*}},G^{\prime}),\\ S^{0}\leftarrow A_{ASA}(S_{code},Alg^{*},G^{\prime}),\end{gathered}$ (3)